# Exploring Neuron Interactions and Emergence in LLMs: From the Multifractal Analysis Perspective

### v4

@@ 6 Discussion and Conclusion
Our novel emergence framework effectively elucidates the interplay between model scale, training steps, and performance, offering a cohesive explanation for the learning evolution in LLMs. Looking ahead, this methodology holds potential not only for enhancing interpretability and assessment, but also for forecasting LLM’s behavior and their subsequent developments. Recent findings by (Bubeck et al., 2023) suggest that current LLMs represent an initial step towards Artificial General Intelligence (AGI). Leveraging our approaches, we might be able to anticipate the conditions necessary for the next leap in model intelligence by our new measurement, potentially marking the threshold for AGI.
@@ A.1.1 Hausdorff Measure
The Hausdorff measure is a key concept in fractal geometry and measure theory, extending the notion of Lebesgue measure to irregular sets, particularly useful for characterizing the size of fractals. It’s named after Felix Hausdorff, a pioneer in set theory and topology.
@@
This dimension is a critical value that separates the scales at which the set $A$ appears “large” from those at which it appears “small” in terms of its $d$-dimensional Hausdorff measure. It provides a precise mathematical way to quantify the notion of dimension for irregular sets, like multifractals, which do not fit neatly into the traditional framework of integer dimensions.
@@ A.1.2 Multifractal Analysis by Box-counting Method
In the domain of network analysis, multifractal analysis (Song et al., 2005, 2007; Furuya & Yakubo, 2011) employs the box-counting method to elucidate the intricate, multifractal structures inherent in complex networks. This analytical approach aligns with the renormalization procedure, essential for understanding multifractality within these systems. The process commences with covering the network using a series of boxes of variable sizes, denoted as $\epsilon$. The count of boxes, $N(\epsilon)$, required to encompass the network at each scale is meticulously recorded. This methodology resonates with the principles of renormalization, highlighting the dynamic interplay between different scales within the network. This approach is akin to assessing the Hausdorff measure A.1.1, a foundational concept in fractal geometry that quantifies the size of a fractal object. Here we provide the process of the calculation of the gerneralized fractal dimension and the multifractal spectrum.
Figure 8: This figure illustrates the renormalization procedure in network theory. It demonstrates the systematic reduction of a network’s scale by grouping nodes within a specified linkage distance, $l_{B}$ into single representative nodes, or ‘boxes’. Each color represents a different box. When two nodes from different original boxes are connected by an edge, the corresponding renormalized nodes are connected as well. This process is iteratively applied, progressively coarsening the network, until it is reduced to a single node. This figure is similar with Figure 1 (a) in (Song et al., 2005).
@@ A.1.3 Power Law in fractal dimension
The definition of fractal dimension is based on unconventional views of scaling and dimension. In conventional geometry, dimension can be defined by intuitive space scaling law. For example, a line with finite length can contain 3 lines with 1/3 its size. But for a square, with a box of side length 1/3 the size of a square, one will find 9 times as many squares as with the original. In general, mathematical definition of such scaling law can be given by
@@
where $N$ stands for the number of measurement units, $\varepsilon$ is the scaling factor. Thus this gives the definition of dimension D:
@@
For lines we can see $N=3$ when $\varepsilon=\frac{1}{3}$. Then we get $D=1$. In the case of a square $D=2$ because $N=9$ when $\varepsilon=\frac{1}{3}$.
@@
where $N_{B}(r_{B})$ is the minimum number of boxes needed to tile a given fractal network with the box radius $r_{B}$. Then we can define the fractal dimension in another way. The mass distribution is defined by the average number of vertices $\langle N(r_{C})\rangle$ within a box with radius $r_{C}$. The mass-radius relation can be given when the relation $M\sim N_{B}(r_{B})\langle N(r_{C})\rangle$ holds for $r_{B}=r_{C}=r$. Then we can characterize the power law between the mass distribution $N(r)$ and the box radius $r$ as
@@
Consider a neuron $v_{i}$ in the layer $l$ of the NIN. The neuron-based fractal dimension $D_{l,i}$ is defined to capture the fractal characteristic based on each neuron. Given the set of distances $\{r_{1},r_{2},...,r_{m}\}$ from $n_{l,i}$ to its neighbors in the next layer $l+1$, the fractal dimension $D_{l,i}$ is calculated as follows:
@@
In practice, to calculate the fractal dimension of a large network we do not need to count all nodes. A practical method is to randomly sample the nodes in the box. Note that when we plot a $\ln{r}-\ln{N}$ graph, the slope is the fractal dimension $D$:
@@ A.2.3 Degree of Emergence
A.3 More Discussion on ‘Self-organization’
@@ A.3.2 Turing pattern
Turing pattern is a famous example of how self-organization occurs in various systems. It was first proposed by Alan Turing when he tried to explain the patterns in morphogenesis. His reaction–diffusion theory has provided insights for understanding pattern formation of various complex objects ranging from chemical molecules to sandpiles.
@@
Here, $D_{i}$ is the diffusion coefficient for object (which could be totally different things in different systems, for example, molecular in chemical process) species $i$ with concentration $C_{i}$, and reactions are described by local non-linear terms included in $\{C_{i}\}$. A stochastic term $\eta(\vec{x},t)$ is also included to describe a general case with stochastic noise: In microscopic system it could result from thermal fluctuations, but for macroscopic ones it could stand for variations in the environment.
@@
  • Morphogens: At least there are two components Activator and Inhibitor. For example, in biological systems they can be signaling molecules.
  • Reaction: For the activator it must be autocatalytic (positive feedback), while the inhibitor could suppress the activator. In short, the reactions (auto-catalysis and cross-catalysis) must exist between them.
  • Diffusion: Both activator and inhibitor diffuse spatially. And the inhibitor diffuses much faster than the activator.
@@
Figure 10: Turing pattern(tur, 2022).
The process of pattern formation is a symmetry-breaking one: at the beginning all spices diffuse in an isotropic way but finally the patterns become like one shown in fig.10. In principle, Turing patterns possess an intrinsic wavelength, which is determined by interactions between molecules and their rates of diffusion. More specifically, it depends only on the ratio of the parameters in the equation, which means it is a scale-less property.
To demonstrate this mathematically, we try to examine the eq.30. We need to find a stable fixed point $\{C_{i}^{*}\}$ as solution. First we apply the linear expansion on the reaction-diffusion equations around the fixed point
@@
Let us come back to talk about the stochastic term in eq.32. For chemical reactions, fluctuations obey typical $\sqrt{N}$ rules for a number of $N$ molecules. We assume the stochastic fluctuations are described by white noise with co-variance
@@
For many variable case in eq.32 the power spectrum of fluctuations in steady state is propotional to $1/\det M(k)$ in eq.40. When $D_{2}/D_{1}>|M_{22}|/M_{11}$, the peak in the power spectrum for noisy fluctuations moves to $k_{m}$.
A.4 Kurtosis ($\beta_{2}$)
Kurtosis is a measure of the ”tailedness” of the probability distribution of a real-valued random variable. It provides insights into the shape of the distribution’s tails and peak. High kurtosis in a data set suggests a distribution with heavy tails and a sharper peak (leptokurtic), while low kurtosis indicates a distribution with lighter tails and a more flattened peak (platykurtic). Kurtosis is often compared to the normal distribution, which has a kurtosis of 3.
The formula for kurtosis is:
$\beta_{2}=E\left[\left(\frac{X-\mu}{\sigma}\right)^{4}\right]$ (50)
where the variables represent the same as in the skewness formula.
A.5 Coefficient of Determination ($R^{2}$)
The coefficient of determination, represented by $R^{2}$, quantifies the extent to which the variance in the dependent variable is predictable from the independent variable(s) in a regression model. It offers a measure of how well observed outcomes are replicated by the model, based on the proportion of total variation of outcomes explained by the model.
The formula to calculate $R^{2}$ is given by:
$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}}$ (51)
$R^{2}$ is interpreted as follows:
Appendix B Model and Dataset Information
Here we will introduce the detailed setting of our experiment mentioned.
B.1 Estimation Method for Network Generation
The time complexity of the algorithm of calculating the shortest paths among each nodes in different layers is $O(kn^{2})$, here $k$ denotes how many layers we crossed and $n$ denotes the number of nodes in each layer. This algorithm calculates the passing closure with a limitation of steps to calculate the shortest paths with a allowance of $k$ layers. Here since the network is a leverage graph, we can reduce the number of nodes we need to compute in each iteration to the number of nodes in each layer. The algorithm is as following:
0:  $Dis$ (A map from node pairs $(u,v)$ to distance value $w$), $Nodes[2...N]$ (An array containing nodes in each layer),$N$ (Number of layers)
1:  for all $x$ in $Nodes[0]$ do
2:     for $l=2$ to $N$ do
3:        for all $y$ in $Nodes[l]$ do
4:           for all $i$ in $Nodes[l-1]$ do
5:              $Dis[x,y]\leftarrow\min\{Dis[x,y],Dis[x,i]+Dis[i,y]\}$
6:           end for
7:        end for
8:     end for
9:  end for
10:  return $Dis$
But the time cost of the calculation of shortest paths in neural networks are unacceptable, because in the large language model, total number of nodes and edges are extremely large. So in real-world situation we need to sample some nodes from each layer to build a reasonable network for analysis. We will explain the sample and shortest path estimation in the appendix.
0:  $Dis$ (A map from node pairs $(u,v)$ to distance value $w$), $Nodes[2...N]$ (An array containing nodes in each layer),$N$ (Number of layers)
1:  $Nodes\leftarrow\text{Sample(}Nodes\text{)}$
2:  for all $x$ in $Nodes[0]$ do
3:     for $l=2$ to $N$ do
4:        for all $y$ in $Nodes[l]$ do
5:           $l\leftarrow 0$
6:           $r\leftarrow\inf$
7:           $hisest\leftarrow 0$
8:           $mid\leftarrow 0$
9:           while $l<r$ do
10:              $mid\leftarrow\frac{l+r}{2}$
11:              $IntNodes\leftarrow\text{Sample}(Nodes[l-1])$
12:              $est\leftarrow\text{ShortestAmong}(IntNodes)$
13:              $hisest,est\leftarrow\min(est,hisest)$
14:              if $est\leq mid$ then
15:                 $r\leftarrow mid$
16:              else
17:                 $l\leftarrow mid$
18:              end if
19:           end while
20:           $dis[x,y]\leftarrow mid$
21:        end for
22:     end for
23:  end for
24:  return $Dis$
Algorithm 5 Calculation of the Partition Function $Z(q)$
0:  $S$ (a dictionary with nodes as keys and lists of distances from the node to its neighbors as values), $q_{\text{list}}$ (a list of distortion factors)
1:  $M_{\text{list}}\leftarrow[]$
2:  $R\leftarrow\{\}$
3:  for all $s$ in $S$ do
4:     $num\leftarrow\text{empty map}$
5:     for all $d$ in $S[s]$ do
6:        $rdist\leftarrow round(d)$
7:        $num[rdist]\leftarrow num[rdist]+1$
8:     end for
9:     $R\leftarrow R\cup\text{keys}(num)$
10:     append $num$ to $M_{\text{list}}$
11:  end for
12:  $r_{\text{sorted}}\leftarrow\text{sort}(R)$
13:  $len_{M}\leftarrow$ length of $M_{\text{list}}$
14:  $len_{r}\leftarrow$ length of $r_{\text{sorted}}$
15:  $W\leftarrow\text{matrix of ones with dimensions }len_{M}\times len_{r}$
16:  for $i=0$ to $len_{M}-1$ do
17:     for $j=0$ to $len_{r}-1$ do
18:        $W[i,j]\leftarrow\sum_{k\leq r_{\text{sorted}}[j]}M_{\text{list}}[i][k]$
19:     end for
20:  end for
21:  $Zq_{\text{list}}\leftarrow[]$
22:  for all $q$ in $q_{\text{list}}$ do
23:     $Zq\leftarrow\sum_{i=0}^{len_{M}}\left(\frac{W[i,:]}{W[i,-1]}\right)^{q}$
24:     append $Zq$ to $Zq_{\text{list}}$
25:  end for
26:  return $(r_{\text{sorted}},Zq_{\text{list}})$
B.2 Model Information
In section 5, we did our experiment on Pythia (Biderman et al., 2023) dataset. The information of this dataset’s model can be found in the table 1.
@@ B.2 Model Information
Table 1: Models in the Pythia suite and select hyperparameters from (Biderman et al., 2023).
And the model being used is based on GPT-NeoX(Andonian et al., 2023). It is an open source library GPT-NeoX developed by EleutherAI. The model checkpoints is saved at initialization, first 1, 2, 4, 8, 16, 32, 64, 128, 256, 512 and every 1000 iterations, making it totally 154 checkpoints.
@@
  • The model uses sparse and dense attention layers in alternation introduced by  Brown et al.. It is used in fully dense layers in the model.
  • The model uses Flash Attention(Dao et al., 2022) during training for improved device throughput.
@@
  • The model uses parallelized attention and feed-forward technique.

### v5

@@ 5 Discussion and conclusion
Our Neuron-based Multifractal Analysis (NeuroMFA) framework addresses a critical gap in quantifying the internal dynamics of large language models (LLMs). By analyzing the multifractal properties of neuron interactions, NeuroMFA offers unique insights into emergent abilities during training. Metrics such as the regularity metric and heterogeneity metric track the evolving complexity and organization within neural networks, providing a cohesive explanation for learning evolution in LLMs. The experimental results demonstrate that emergence in LLMs can be explained by modeling and analyzing the structural and self-organization properties among neurons, aligning with theories of brain function. While accessing human brain neuronal parameters remains challenging, LLMs provide a transparent and accessible platform for studying self-organization phenomena, making this research particularly promising. In future applications, NeuroMFA holds potential not only for enhancing interpretability and assessment but also for forecasting LLMs’ behavior and guiding their development. Moreover, by linking structure to the capability of large artificial models, NeuroMFA enables a deeper exploration of potential connections between artificial and biological intelligence.
@@ A.1.1 Hausdorff measure
The Hausdorff measure is a key concept in fractal geometry and measure theory, extending the notion of Lebesgue measure to irregular sets, and is particularly useful for characterizing the size of fractals. It’s named after Felix Hausdorff, a pioneer in set theory and topology.
@@
This dimension is a critical value that separates the scales at which the set $A$ appears “large" from those at which it appears “small" in terms of its $d$-dimensional Hausdorff measure. It provides a precise mathematical way to quantify the notion of dimension for irregular sets, like multifractals, which do not fit neatly into the traditional framework of integer dimensions.
@@ A.1.2 Multifractal analysis by box-counting method
In the domain of network analysis, multifractal analysis (Song et al., 2005; 2007; Furuya & Yakubo, 2011) employs the box-counting method to elucidate the intricate, multifractal structures inherent in complex networks. This analytical approach aligns with the renormalization procedure, essential for understanding multifractality within these systems. The process commences with covering the network using a series of boxes of variable sizes, denoted as $\epsilon$. The count of boxes, $N(\epsilon)$, required to encompass the network at each scale is meticulously recorded. This methodology resonates with the principles of renormalization, highlighting the dynamic interplay between different scales within the network. This approach is akin to assessing the Hausdorff measure A.1.1, a foundational concept in fractal geometry that quantifies the size of a fractal object. Here we provide the process of the calculation of the generalized fractal dimension and the multifractal spectrum.
Figure 9: This figure illustrates the renormalization procedure in network theory. It demonstrates the systematic reduction of a network’s scale by grouping nodes within a specified linkage distance, $l_{B}$ into single representative nodes, or ‘boxes’. Each color represents a different box. When two nodes from different original boxes are connected by an edge, the corresponding renormalized nodes are connected as well. This process is iteratively applied, progressively coarsening the network, until it is reduced to a single node. This figure is similar to Fig. 1 (a) in Song et al. (2005).
@@ A.1.3 Power law in fractal dimension
The definition of fractal dimension is based on unconventional views of scaling and dimension. In conventional geometry, dimension can be defined by the intuitive space scaling law. For example, a line with finite length can contain 3 lines with 1/3 its size. But for a square, with a box of side length 1/3 the size of a square, one will find 9 times as many squares as with the original. In general, mathematical definition of such a scaling law can be given by
@@
where $N$ stands for the number of measurement units, $\varepsilon$ is the scaling factor. Thus, this gives the definition of dimension D:
@@
For lines, we can see $N=3$ when $\varepsilon=\frac{1}{3}$. Then we get $D=1$. In the case of a square $D=2$ because $N=9$ when $\varepsilon=\frac{1}{3}$.
@@
where $N_{B}(r_{B})$ is the minimum number of boxes needed to tile a given fractal network with the box radius $r_{B}$. Then we can define the fractal dimension in another way. The mass distribution is defined by the average number of vertices $\langle N(r_{C})\rangle$ within a box with radius $r_{C}$. The mass-radius relation can be given when the relation $M\sim N_{B}(r_{B})\langle N(r_{C})\rangle$ holds for $r_{B}=r_{C}=r$. we can then characterize the power law between the mass distribution $N(r)$ and the box radius $r$ as
@@
Consider a neuron $v_{i}$ in layer $l$ of the NIN. The neuron-based fractal dimension $D_{l,i}$ is defined to capture the fractal characteristic based on each neuron. Given the set of distances $\{r_{1},r_{2},...,r_{m}\}$ from $n_{l,i}$ to its neighbors in the next layer $l+1$, the fractal dimension $D_{l,i}$ is calculated as follows:
@@
In practice, to calculate the fractal dimension of a large network we do not need to count all nodes. A practical method is to randomly sample the nodes in the box. Note that, when we plot a $\ln{r}-\ln{N}$ graph, the slope is the fractal dimension $D$:
@@ A.2.3 Degree of emergence
A.2.4 Illustration examples
We provide the concrete examples to illustrate the multifractality and monofractality in Fig. 11. Cantor sets are classical fractal structures created by iteratively removing the middle part of a line segment. Fig. 11 (a) shows the patterns with self-similar statistical properties characterized by a single fractal dimension across all scales while Fig. 11 (b) shows the patterns characterized by different fractal dimensions across multiple scales. Here, the mass probability $p_{i}$ is defined as the ratio of the number of segments in each subinterval to the total number of segments. By applying the multifractal analysis, as shown in Fig. 11 (c), the spectrum of a monofractal Cantor set results in a single point, indicating uniform scale characteristics. In contrast, the multifractal Cantor set’s spectrum forms a bell-shaped curve, reflecting the diversity of scale characteristics.
In our NeuroMFA, the mass probability is represented as the ratio of nodes in a box to the total number of nodes, with the observation scale being the radius of the box (see Figs. 11 (f-g)). By measuring and distorting the mass distribution in different boxes, we can capture the node-based multifractal characteristics in the network. In Figs. 11 (d-e), we provide an example of a Watts-Strogatz (WS) network model transitioning from a regular network to a completely random network. A completely regular network exhibits monofractal properties, with the multifractal spectrum being a single point. As edges are randomly rewired, the network structure’s diversity increases, leading to an increased spectrum width. Simultaneously, the network structure becomes more irregular, indicated by a right-shifted spectrum.
Figure 11: Concrete examples of monofractal and multifractal. Fig. (a) illustrates self-similar patterns with a single fractal dimension across all scales, while Fig. (b) depicts patterns with varying fractal dimensions across multiple scales. Fig. (c) shows the monofractal Cantor set yielding a single-point spectrum, contrasted by the bell-shaped multifractal spectrum for a multifractal Cantor set. In Figs. (d-e), a Watts-Strogatz (WS) network model transitions from a regular to a random network, where rewiring edges increases spectrum width and irregularity, shifting the multifractal spectrum. Figs. (f-g) demonstrate the node-based multifractal analysis in a network, using box-counting techniques.
Figure 12: Illustration of the proposed metrics. (a-c) Experiment 1 illustrates the transformation from a homogeneous to a more heterogeneous network, as shown by (g) broadened spectral width. (d-f) Experiment 2 demonstrates increasing regularity, starting from random configurations and progressing through multiple Weighted Preferential Attachment Mechanism (wPAM) iterations, reflected in (h) the shift and decrease in $\alpha_{0}$.
To intuitively illustrate the metrics, we apply NeuroMFA to toy examples. These metrics provide quantitative insights into the heterogeneity and regularity of neuronal networks. Using simplified models consisting of three layers with 32 nodes each, we demonstrate how network transformations influence these metrics through two experiments:
  • Experiment 1: Focused on the effects of increasing network heterogeneity. Starting with a homogeneous network, we systematically varied node characteristics to simulate increased complexity.
  • Experiment 2: Explored the introduction of regularity into a network using Algorithm 1 . We began with a randomly structured network and used the wPAM to gradually increase regularity.
The results are visualized in Fig. 12, showing distinct phases of the experiments and how NeuroMFA metrics quantify changes in network heterogeneity and regularity.
Algorithm 1 Multiple Rounds of Weight Adjustment
1:$N$ (number of iterations)
2:for $n=1$ to $N$ do
3:     for $i=1$ to 128 do
4:         for $j=1$ to 32 do
5:              if $W_{ij}<\text{median}(W)$ then
6:                  $W_{ij}\leftarrow W_{ij}\cdot\delta$triangleright Decrease weight
7:              else
8:                  $W_{ij}\leftarrow W_{ij}\cdot\iota$triangleright Increase weight
9:              end if
10:         end for
11:     end for
12:end for
A.3 Discussion on self-organization
@@ A.3.2 Turing pattern
Turing pattern is a well-known example of how self-organization occurs in various systems. It was first proposed by Alan Turing when he tried to explain the patterns in morphogenesis. His reaction–diffusion theory has provided insights for understanding pattern formation of various complex objects ranging from chemical molecules to sandpiles.
@@
Here, $D_{i}$ is the diffusion coefficient for object species $i$ (which could be totally different things in different systems, for example, molecular in chemical process) with concentration $C_{i}$, and reactions are described by local non-linear terms included in $\{C_{i}\}$. A stochastic term $\eta(\vec{x},t)$ is also included to describe the general case with stochastic noise: In microscopic system it could result from thermal fluctuations, whereas for macroscopic ones it could stand for variations in the environment.
@@
  • Morphogens: There must be at least two components Activator and Inhibitor. For example, in biological systems, they can be signaling molecules.
  • Reaction: For the activator it must be autocatalytic (positive feedback), while the inhibitor could suppress the activator. In short, the reactions (auto-catalysis and cross-catalysis) must occur between them.
  • Diffusion: Both activator and inhibitor diffuse spatially. And the inhibitor diffuse much faster than the activator.
@@
Figure 13: Turing pattern(Turing, 2022).
The process of pattern formation is a symmetry-breaking one: at the beginning all spices diffuse in an isotropic way but finally the patterns become like the one shown in Fig.13. In principle, Turing patterns possess an intrinsic wavelength, which is determined by interactions between molecules and their rates of diffusion. More specifically, it depends only on the ratio of the parameters in the equation, which means it is a scale-independent property.
To demonstrate this mathematically, we try to examine the Eq.30. We need to find a stable fixed point $\{C_{i}^{*}\}$ as solution. First, we apply the linear expansion on the reaction-diffusion equations around the fixed point
@@
Let us come back to talk about the stochastic term in Eq.32. For chemical reactions, fluctuations obey typical $\sqrt{N}$ rules for a number of $N$ molecules. We assume the stochastic fluctuations are described. by white noise with co-variance
@@
For many variable case in Eq.32 the power spectrum of fluctuations in steady state is proportional to $1/\det M(k)$ in Eq.38. When $D_{2}/D_{1}>|M_{22}|/M_{11}$, the peak in the power spectrum for noisy fluctuations shifts to $k_{m}$.
A.3.3 Emergence and self-organization in Large Models
“Emergence refers to the arising of novel and coherent structures, patterns, and properties during the process of self-organization” (Goldstein, 1999). Emergence can be described as the formation of new properties - “something new appears” - in the evolution of a collective system that cannot be fully captured by its dynamical equations. These emergent properties arise from local interactions and can be harnessed by the system to develop additional functionalities or capabilities (Crutchfield, 1994). In the context of large models, emergence is the complex behavior through which new learning capabilities are acquired by the LLM from the collective interactions among numerous simple elements (i.e., the neurons following consistent rules within the neural network).
“Self-organization is the process of forming patterns and achieving overall order through internal local interactions” (Correia, 2006; Ay et al., 2007). In the human brain, intelligence and complex behaviors arise from the collective behavior of neurons, with the brain operating as a self-organizing system (Kelso, 1995; Singer, 1986; Pribram, 2018). In the context of large models, self-organization refers to the process where the interactions of artificial neurons during training evolve from initial randomness, developing new patterns and orderly structures, thus leading to emergent abilities in large models.
Appendix B Implementation details
B.1 Code availability
The source code is available at https://anonymous.4open.science/r/Neuron_LLM-4DC8. Here we provide the hyper-parameters we used to produce our experiments in Table 2.
B.2 Evaluation LLMs
We experiment on Pythia Biderman et al. (2023) dataset and models. The detailed information can be found in the table 1.
@@ B.2 Evaluation LLMs
Table 1: Models in the Pythia suite and select hyperparameters from Biderman et al. (2023).
And the model being used is based on GPT-NeoXAndonian et al. (2023). It is an open source library GPT-NeoX developed by EleutherAI. The model checkpoints are saved at initialization, at the first 1, 2, 4, 8, 16, 32, 64, 128, 256, 512 iterations, and then every 1000 iterations, making it a total of 154 checkpoints.
@@
  • The model uses sparse and dense attention layers in alternation introduced by  Brown et al.. It is used in the fully dense layers of the model.
  • The model uses Flash AttentionDao et al. (2022) during training for improved device throughput.
@@
  • The model uses a parallelized attention and feed-forward technique.